OpenStudy (anonymous):

Solve the system by elimination. -2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5

OpenStudy (campbell_st):

well you need to reduce it to 2 equations in 2 unknowns... so to eliminate x add the 1st and 3rd equation which gives 5y + 6z = 5 and the 2nd and 3rd equations 3y + 4z = 2 so you are now looking at 5y + 6z = 5 3y + 4z = 2 now choose a letter to eliminate.... you'll find the answers will be mixed numbers...

OpenStudy (anonymous):

I don't get how to explain the answer if I put it into a calculator.

OpenStudy (campbell_st):

well its a bit early for a calculator... unless it is an Nspire that will solve this so using the equations 5y + 6z = 5 (4) 2y + 4x = 2 (5) take the 5th equation and make y the subject so 2y = 2 - 4z or y = 1 - 2z substitute it into the 4th equation 5(1 - 2z) + 6z = 5 so 5 - 10z + 6x = 5 which means -4z + 5 = 5 or z = 0 substitute this into the 4th equation 5y + 6(0) = 5 so y = 1 now substiute z = 0 and y = 1 into the 1st equation and you get -2x + 2(1) + 3(0) = 0 or -2x + 2 = 0 solve for x then you'll have the solutions x = , y = 1 and z = 0